# Recursive decoding of projective Reed-Muller codes

### v1

@@ Abstract.
We give a recursive decoding algorithm for projective Reed-Muller codes making use of a decoder for affine Reed-Muller codes. We determine the number of errors that can be corrected in this way, which is the current highest for decoders of projective Reed-Muller codes. We show when we can decode up to the error correction capability of these codes, and we compute the order of complexity of the algorithm, which is given by that of the chosen decoder for affine Reed-Muller codes.
@@ 1. Introduction
Projective Reed-Muller (PRM) codes are a family of evaluation codes introduced in [23]. They are obtained by evaluating multivariate homogeneous polynomials in the projective space, and their basic parameters were completely determined in [34, 12]. Their affine counterpart, affine Reed-Muller (RM) codes, has been widely studied. We know the basic parameters of RM codes [20, 7], but also many additional properties [15, 22, 36, 3, 2, 1, 21]. There has been some recent work in the direction of closing the gap in knowledge between RM and PRM codes [5, 32, 27, 28, 12, 14, 13, 19, 33]. However, one of the main aspects required for a family of codes to be useful in practice is to have an efficient decoding algorithm. RM codes admit several efficient decoding algorithms, most of them derived from the fact that they can be seen as algebraic geometry codes [29, 30, 31, 10, 11]. For PRM codes, the only specific decoder was given in [26]. This decoder is efficient in terms of complexity, but does not decode up to the error correction capability of PRM codes.
@@
In [32], a recursive construction for PRM codes is given. This recursive construction is similar to a $(u,u+v)$ construction, which is known to have an efficient decoding algorithm [16, 17]. Moreover, affine RM codes also admit a recursive construction which has been used to obtain decoding algorithms [37, 8]. In this paper we propose an algorithm based on the recursive construction from [32] to decode PRM codes.
@@
In Section 2 we provide the necessary background. In this section we also introduce a particular ordering for the points of the projective space, which is a key element for the recursive structure of PRM codes. Moreover, in Equation (3) we define an integer $\eta_{d}(m)$, which will determine the error correction capability of our algorithm, and we study some of its properties, e.g., when we have the equality between $\eta_{d}(m)$ and the minimum distance of PRM codes. In Section 3 we study how to decode PRM codes recursively using decoders for RM codes. First, in Subsection 3.1 we show how this idea gives a decoding algorithm for projective Reed-Solomon (PRS) which decodes up to their error correction capability, using a decoder for affine Reed-Solomon (RS) codes. Then, in Subsection 3.2 we explain how the algorithm works for the case of PRM codes over the projective plane. This already features most of the aspects of the general recursive algorithm, but it is easier to grasp, and serves as an intermediate step between PRS codes and the general case of PRM codes. We also provide examples to help understand some of the underlying problems we have to solve to successfully decode PRM codes. In Subsection 3.3, we give the recursive decoding algorithm, Algorithm 1, and we prove that it can decode any number of errors lower than $\eta_{d}(m)/2$. By the definition of $\eta_{d}(m)$, this means that we can always decode at least the same number of errors as the algorithm presented in [26], but in general we will be able to correct more. In fact, due to Lemma 2.9, we can study when Algorithm 1 can decode up to the error correction capability of PRM codes. This is done in Section 4, where we also consider an improved algorithm, Algorithm 2, which can additionally decode some particular error patterns even if the weight of the error exceeds the aforementioned error correction capability of Algorithm 1. In Section 5 we compute the order of complexity of Algorithms 1 and 2, which is the same as the complexity order of the chosen decoder for RM codes. Finally, in Section 6 we mention some future avenues of research.
@@ 2. Preliminaries
Let $\mathbb{F}_{q}$ be the finite field of $q$ and let $m$ be a positive integer. We consider the projective space $\mathbb{P}^{m}$ over $\mathbb{F}_{q}$, and we denote by $p_{j}$ the number of points in $\mathbb{P}^{j}$, i.e., $p_{j}=\left\lvert\mathbb{P}^{j}\right\rvert=\frac{q^{j+1}-1}{q-1}$. Throughout this work we will fix the standard representatives for $\mathbb{P}^{m}$, which are the representatives with the leftmost nonzero coordinate equal to 1. The set $P^{m}$ whose elements are these representatives can be regarded as a subset of the affine space:
@@
were we use the convention $P^{0}:=\{1\}$. In what follows, we will assume a specific ordering of the points in $P^{m}$. First, let $\xi\in\mathbb{F}_{q}$ be a primitive element. We order the elements of $\mathbb{F}_{q}$ in terms of the powers of $\xi$, that is
@@
Indeed, given a point $Q$ in ${\mathbb{F}}_{q}^{m}\setminus\{(0,\dots,0\}$, its leftmost nonzero coordinate is equal to $\xi^{r}$ for some $20\leq r\leq q-$ 2, which implies that $Q\in\xi^{r}\cdot P^{m-1}$. In particular, we have
@@
Let $d$ be a positive integer, and let $\mathbb{F}_{q}[x_{0},\dots,x_{m}]_{d}$ be the set of homogeneous polynomials of degree $d$. Then $\operatorname{PRM}_{d}(q,j):=\operatorname{ev}^{\mathbb{P}}_{{j}}(\mathbb{F}_{q}[x_{m-j},\dots,x_{m}]_{d})$ is the PRM code of degree $d$ over $P^{j}$. We may use the notation $\operatorname{PRM}_{d}(j)$ if there is no confusion about the field. We have introduced the definition of $\operatorname{ev}^{\mathbb{P}}_{{j}}$ and $\operatorname{PRM}_{d}(j)$ because it will be useful for the decoding algorithm, but we will mainly focus in what follows on $\operatorname{PRM}_{d}(m)$. For $m=1$, we obtain PRS codes (sometimes called doubly extended Reed-Solomon codes), which are MDS codes with parameters $[q+1,d+1,q-d+1]$. We use the notation $\operatorname{PRS}_{d}=\operatorname{PRM}_{d}(1)$. The following result from [34] gives the basic parameters of these codes (also see [12, 35] for the minimum distance).
@@
$\operatorname{ev}^{\mathbb{A}}_{{j}}:\mathbb{F}_{q}[x_{m-j+1},\dots,x_{m}]\rightarrow\mathbb{F}_{q}^{q^{j}},\>\>f\mapsto\left(f(Q_{1}),\dots,f(Q_{n})\right)_{Q_{i}\in\mathbb{F}_{q}^{j}}.$
If we denote by $\mathbb{F}_{q}[x_{m-j+1},\dots,x_{m}]_{\leq d}$ the polynomials of degree less than or equal to $d$, we have $\operatorname{RM}_{d}(j):=\operatorname{ev}^{\mathbb{A}}_{{j}}(\mathbb{F}_{q}[x_{m-j+1},\dots,x_{m}]_{\leq d})$. As with PRM codes, we will mainly consider the code $\operatorname{RM}_{d}(m)$. For $m=1$ we obtain Reed-Solomon codes, for which we use the notation $\operatorname{RS}_{d}=\operatorname{RM}_{d}(1)$. From [7, 20] we have the following result about the basic parameters of RM codes.
@@
[…]
Then every class in $\mathbb{F}_{q}[x_{1},\dots,x_{m}]/I({\mathbb{A}}^{m})$ has one representative (and only one) in $A(m)$. Thus, whenever we consider a polynomial $f\in\mathbb{F}_{q}[x_{1},\dots,x_{m}]$, we denote by $f\bmod I({\mathbb{A}}^{m})$ its representative in $\mathbb{F}_{q}[x_{1},\dots,x_{m}]/I({\mathbb{A}}^{m})$ whose monomials are contained in $A(m)$. We say that $f$ is reduced modulo $I({\mathbb{A}}^{m})$ if it is expressed in terms of the monomials in $A(m)$. We can now give the following result, which is related to Remark 2.3 and will be used later.
@@
If $x_{0}\mid x^{\alpha}$ and $x_{0}\nmid x^{\beta}$ (recall $a\geq b$), we have that $x^{\beta}(1,x_{1},\dots,x_{m})=x^{\beta}$ is of degree $d$, and $x^{\alpha}(1,x_{1},\dots,x_{m})$ is of degree $<d$. Thus, in order to have
@@
the reduction of $x^{\beta}$ modulo modulo $I({\mathbb{A}}^{m})$ has to be different from $x^{\beta}$, which implies that the degree of at least one variable in $x^{\beta}$ is higher than $q-1$. Then
@@
Note that $h_{d}^{j}(A(j))\subset M_{d}(j)$. Now we introduce additional notation related to homogenizations which will be needed for Algorithm 1.
@@
Let $1\leq d\leq m(q-1)$. Let $x^{\alpha}\in\mathbb{F}_{q}[x_{1},\dots,x_{m}]$ be a monomial. We say that $x^{\alpha}$ is bad if $0<\left\lvert\alpha\right\rvert<d$ and $\left\lvert\alpha\right\rvert\equiv d\bmod q-1$. We say that $x^{\alpha}$ is good if it is not bad. Given a polynomial $f\in\mathbb{F}_{q}[x_{1},\dots,x_{m}]_{\leq d}$, we denote by ${f_{bad}}$ the sum of the terms of $f$ corresponding to bad monomials. Let ${f_{good}}=f-{f_{bad}}$. We denote by $({f_{good}})_{d}$ and $({f_{good}})_{\leq d-1}$ the sum of the terms of degree $d$ and degree less than or equal to $d-1$ of ${f_{good}}$, respectively.
@@
As a particular case of the implications of the recursive construction for the GHWs of PRM codes given in [32, Thm. 7], this construction gives a bound $\eta_{d}(m)$ for the minimum distance of PRM codes which will determine the error correction capability of the recursive decoding algorithm that we will develop in what follows. Let $d-1=\nu(q-1)+\mu$, with $10\leq\mu<q-$ 1. We denote
(3) ${\eta_{d}(m)}:=\sum_{i=0}^{m-\nu-1}\operatorname{wt}(\operatorname{RM}_{d}(m-i))+1$
@@
Let $d-1=\nu(q-1)+\mu$, with $10\leq\mu\leq q-$ 1. Note that $d-(q-1)-1=(\nu-1)(q-1)+\mu$, and, by Lemma 2.9, we have
@@
As we mentioned above, the lower bound ${\eta_{d}(m)}$ for $\operatorname{wt}(\operatorname{PRM}_{d}(m))$ is motivated by the recursive construction from [32]. Indeed, from [32, Thm. 7] we have
@@
For $d>(m-1)(q-1)$, we have $\operatorname{wt}(\operatorname{RM}_{d-1}(m))\leq q\cdot\operatorname{wt}(\operatorname{PRM}_{d}(m-1))=q$. Thus, for $d\leq m(q-1)$ we obtain
@@
$\operatorname{ev}^{\mathbb{A}}_{{m}}(f(1,x_{1},\dots,x_{m}))=\operatorname{ev}^{\mathbb{P}}_{{m}}(f(x_{0},x_{1},\dots,x_{m}))\mid_{\{1\}\times\mathbb{F}_{q}^{m}}=(u+v_{\xi,d}),$
where, for $w\in\mathbb{F}_{q}^{q^{m}}$, we denote $w\mid_{\{1\}\times\mathbb{F}_{q}^{m}}:=(w_{Q})_{Q\in\{1\}\times\mathbb{F}_{q}^{m}}$.
@@ 3. Recursive decoding algorithm
Fix $m>0$. In this section we show a decoding algorithm, Algorithm 1, to decode $\operatorname{PRM}_{d}(m)$ if the number of errors is lower than $t_{0}:=\eta_{d}(m)/2$. We denote the corresponding decoder $D^{\mathbb{P}}_{d}(m)$. This algorithm relies on a choice of a decoding algorithm $D^{\mathbb{A}}_{d}(j)$ for RM codes, for each $j\leq m$. To the knowledge of the author, the most efficient decoding algorithm is obtained by considering the Berlekamp-Massey-Sakata algorithm [29, 30] together with the majority voting algorithm [10, 11] to decode RM codes up to their error correction capability [31], or even more depending on the type of errors [6]. A simpler approach could be to just consider majority voting decoding as in [18, 25, 11], which also decodes up to the error correction capability of RM codes, but with higher complexity. With respect to PRM codes, the only specific decoder known is given in [26], in which the authors consider a decomposition of the projective space into affine spaces, while our approach relies on the recursive construction from [32].
@@
As a convention, if at some point we use a decoder whose error correction capability is 00, we assume it returns the received vector if it belongs to the corresponding code, and an error otherwise. With respect to the terminology, by the error correction capability of an $[n,k,\delta]$ code $C$ we mean $\left\lfloor\frac{\delta-1}{2}\right\rfloor$, since it can correct any number of errors lower than or equal to $\left\lfloor\frac{\delta-1}{2}\right\rfloor$, or, equivalently, any number of errors lower than $\delta/2$. For completeness, we start by briefly covering the case of PRS codes, then we show how our decoder works for $\operatorname{PRM}_{d}(2)$ to illustrate the main ideas of Algorithm 1, and then we prove that the algorithm works in the general case of $\operatorname{PRM}_{d}(m)$.

### v2

@@ Abstract.
We give a recursive decoding algorithm for projective Reed-Muller codes making use of a decoder for affine Reed-Muller codes. We determine the number of errors that can be corrected in this way, which is the current highest for decoders of projective Reed-Muller codes. We determine the degrees for which we can decode up to the error correction capability of these codes, and we compute the order of complexity of the algorithm, given by that of the chosen decoder for affine Reed-Muller codes. We also show how the decoder behaves when increasing the number of variables and the size of the field.
@@ 1. Introduction
Projective Reed-Muller (PRM) codes are a family of evaluation codes introduced in [25]. They are obtained by evaluating multivariate homogeneous polynomials in the projective space, and their basic parameters were completely determined in [36, 13]. Their affine counterpart, affine Reed-Muller (RM) codes, have been widely studied. We know the basic parameters of RM codes [21, 8], but also many additional properties [16, 24, 38, 3, 2, 1, 22]. There has been some recent work in the direction of closing the gap in knowledge between RM and PRM codes [5, 34, 29, 30, 13, 15, 14, 20, 35]. However, one of the main aspects required for a family of codes to be useful in practice is to have an efficient decoding algorithm. RM codes admit several efficient decoding algorithms, most of them derived from the fact that they can be seen as algebraic geometry codes [31, 32, 33, 11, 12]. For PRM codes, the only specific decoder was given in [28]. This decoder is efficient in terms of complexity, but does not decode up to the error correction capability of PRM codes.
@@
In [34], a recursive construction for PRM codes is given. This recursive construction is similar to a $(u,u+v)$ construction, which is known to have an efficient decoding algorithm [17, 18]. Moreover, affine RM codes also admit a recursive construction which has been used to obtain decoding algorithms [39, 9]. In this paper, we propose an algorithm based on the recursive construction from [34] to decode PRM codes.
@@
In Section 2 we provide the necessary background. In this section, we also introduce a particular ordering for the points of the projective space, which is a key element for the recursive structure of PRM codes. We define an integer $\eta_{d}(m)$, which will determine the error correction capability of our algorithm, and we study some of its properties. Since in general $\eta_{d}(m)$ is only a lower bound for the minimum distance of the corresponding PRM code, the decoder may not decode up to the error correction capability of the code for all possible degrees, and we determine the degrees for which it does. In Section 3, we study how to decode PRM codes recursively using decoders for RM codes. First, in Subsection 3.1 we show how this idea gives a decoding algorithm for projective Reed-Solomon (PRS) which decodes up to their error correction capability, using a decoder for affine Reed-Solomon (RS) codes. Then, in Subsection 3.2 we explain how the algorithm works for the case of PRM codes over the projective plane. This already features most of the aspects of the general recursive algorithm, but it is easier to grasp and serves as an intermediate step between PRS codes and the general case of PRM codes. We also provide examples to help understand some of the underlying problems we have to solve to successfully decode PRM codes. In Subsection 3.3, we give the recursive decoding algorithm, Algorithm 1, and we prove that it can decode any number of errors lower than $\eta_{d}(m)/2$. By the definition of $\eta_{d}(m)$, this means that we can always decode at least the same number of errors as the algorithm presented in [28], but in general, we will be able to correct more. In fact, we study for which degrees Algorithm 1 can decode up to the error correction capability of PRM codes. This is done in Section 4, where we also consider an improved algorithm, Algorithm 2, which can additionally decode some particular error patterns even if the weight of the error exceeds the aforementioned error correction capability of Algorithm 1. In Section 5 we compute the order of complexity of Algorithms 1 and 2, which is the same as the complexity order of the chosen decoder for RM codes. Finally, in Section 6 we mention some future avenues of research.
@@ 2. Preliminaries
Let $\mathbb{F}_{q}$ be the finite field of $q$ and let $m$ be a positive integer. We consider the projective space $\mathbb{P}^{m}$ over $\mathbb{F}_{q}$, and we denote by $p_{j}$ the number of points in $\mathbb{P}^{j}$, i.e., $p_{j}=\left\lvert\mathbb{P}^{j}\right\rvert=\frac{q^{j+1}-1}{q-1}$. Throughout this work, we will fix the standard representatives for $\mathbb{P}^{m}$, which are the representatives with the leftmost nonzero coordinate equal to 1. The set $P^{m}$ whose elements are these representatives can be regarded as a subset of the affine space:
@@
where we use the convention $P^{0}:=\{1\}$. In what follows, we will assume a specific ordering of the points in $P^{m}$. First, let $\xi\in\mathbb{F}_{q}$ be a primitive element. We order the elements of $\mathbb{F}_{q}$ in terms of the powers of $\xi$, that is
@@
Indeed, given a point $Q$ in ${\mathbb{F}}_{q}^{m}\setminus\{(0,\dots,0\}$, its leftmost nonzero coordinate is equal to $\xi^{r}$ for some $0\leq r\leq q-2$, which implies that $Q\in\xi^{r}\cdot P^{m-1}$. In particular, we have
@@
Let $d$ be a positive integer, and let $\mathbb{F}_{q}[x_{0},\dots,x_{m}]_{d}$ be the set of homogeneous polynomials of degree $d$. Then $\operatorname{PRM}_{d}(q,j):=\operatorname{ev}^{\mathbb{P}}_{{j}}(\mathbb{F}_{q}[x_{m-j},\dots,x_{m}]_{d})$ is the PRM code of degree $d$ over $P^{j}$. We may use the notation $\operatorname{PRM}_{d}(j)$ if there is no confusion about the field. We have introduced the definition of $\operatorname{ev}^{\mathbb{P}}_{{j}}$ and $\operatorname{PRM}_{d}(j)$ because it will be useful for the decoding algorithm, but we will mainly focus on $\operatorname{PRM}_{d}(m)$ in what follows. For $m=1$, we obtain PRS codes (sometimes called doubly extended Reed-Solomon codes), which are MDS codes with parameters [q+1,d+1,qd+1][q+1,d+1,q-d+1]. We use the notation $\operatorname{PRS}_{d}=\operatorname{PRM}_{d}(1)$. The following result from [36] gives the basic parameters of these codes (also see [13, 37] for the minimum distance).
@@
$\operatorname{ev}^{\mathbb{A}}_{{j}}:\mathbb{F}_{q}[x_{m-j+1},\dots,x_{m}]\rightarrow\mathbb{F}_{q}^{q^{j}},\>\>f\mapsto\left(f(Q_{1}),\dots,f(Q_{q^{j}})\right)_{Q_{i}\in\mathbb{F}_{q}^{j}}.$
If we denote by $\mathbb{F}_{q}[x_{m-j+1},\dots,x_{m}]_{\leq d}$ the polynomials of degree less than or equal to $d$, we have $\operatorname{RM}_{d}(j):=\operatorname{ev}^{\mathbb{A}}_{{j}}(\mathbb{F}_{q}[x_{m-j+1},\dots,x_{m}]_{\leq d})$. As with PRM codes, we will mainly consider the code $\operatorname{RM}_{d}(m)$. For $m=1$ we obtain Reed-Solomon codes of length $q$, for which we use the notation $\operatorname{RS}_{d}=\operatorname{RM}_{d}(1)$. From [8, 21] we have the following result about the basic parameters of RM codes.
@@
$A(j):=\{x^{\alpha}\in\mathbb{F}_{q}[x_{m-j+1},\dots,x_{m}]:0\leq\alpha_{i}\leq q-1,\;1\leq i\leq m\}.$
Then every class in $\mathbb{F}_{q}[x_{1},\dots,x_{m}]/I({\mathbb{A}}^{m})$ has one representative (and only one) in $A(m)$. Thus, whenever we consider a polynomial $f\in\mathbb{F}_{q}[x_{1},\dots,x_{m}]$, we denote by $f\bmod I({\mathbb{A}}^{m})$ its representative in $\mathbb{F}_{q}[x_{1},\dots,x_{m}]/I({\mathbb{A}}^{m})$ whose monomials are contained in $A(m)$. We say that $f$ is reduced modulo $I({\mathbb{A}}^{m})$ if it is expressed in terms of the monomials in $A(m)$. We can now give the following result, which is related to Remark 2.3 and will be used later. We use the following notation: given a monomial $x^{\alpha}=x_{0}^{\alpha_{0}}\cdots x_{m}^{\alpha_{m}}\in\mathbb{F}_{q}[x_{0},\dots,x_{m}]$, we denote by $x^{\alpha}(1,x_{1},\dots,x_{m})$ the monomial $x_{1}^{\alpha_{1}}\cdots x_{m}^{\alpha_{m}}$.
@@
If $x_{0}\mid x^{\alpha}$ and $x_{0}\nmid x^{\beta}$ (recall $a\geq b$), we have that $x^{\beta}(1,x_{1},\dots,x_{m})=x^{\beta}$ is of degree $d$, and $x^{\alpha}(1,x_{1},\dots,x_{m})$ is of degree <d<d. Thus, in order to have
@@
the reduction of $x^{\beta}$ modulo $I({\mathbb{A}}^{m})$ has to be different from $x^{\beta}$, which implies that the degree of at least one variable in $x^{\beta}$ is higher than q1q-1. Then
@@
Note that $h_{d}^{j}(A(j))\subset M_{d}(j)$. Now we introduce additional notation related to homogenizations, which will be needed for Algorithm 1.
@@
Let $1\leq d\leq m(q-1)$. Let $x^{\alpha}\in\mathbb{F}_{q}[x_{1},\dots,x_{m}]$ be a monomial. We say that $x^{\alpha}$ is bad if $0<\left\lvert\alpha\right\rvert<d$ and $\left\lvert\alpha\right\rvert\equiv d\bmod q-1$. We say that $x^{\alpha}$ is good if it is not bad. Given a polynomial $f\in\mathbb{F}_{q}[x_{1},\dots,x_{m}]_{\leq d}$, we denote by $f_{bad}$} the sum of the terms of $f$ corresponding to bad monomials. Let ${f_{good}}=f-{f_{bad}}$. We denote by $({f_{good}})_{d}$ and $({f_{good}})_{\leq d-1}$ the sum of the terms of degree $d$ and degree less than or equal to d1d-1 of $f_{good}$}, respectively.
@@
As a particular case of the implications of the recursive construction for the generalized Hamming weights of PRM codes given in [34, Thm. 7], this construction gives a bound $\eta_{d}(m)$ for the minimum distance of PRM codes, which will determine the error correction capability of the recursive decoding algorithm that we will develop in what follows. Let $d-1=\nu(q-1)+\mu$, with $0\leq\mu<q-1$. We denote
(3) ${\eta_{d}(m)}=\sum_{i=0}^{m-\nu-1}\operatorname{wt}(\operatorname{RM}_{d}(m-i))+1$
@@
Let $d-1=\nu(q-1)+\mu$, with $0\leq\mu\leq q-1$. Note that $d-(q-1)-1=(\nu-1)(q-1)+\mu$, and, by Lemma 2.9, we have
@@
As we mentioned above, the lower bound $\eta_{d}(m)$} for $\operatorname{wt}(\operatorname{PRM}_{d}(m))$ is motivated by the recursive construction from [34]. Indeed, from [34, Thm. 7] we have
@@
For d>(m1)(q1)d>(m-1)(q-1), we have $\operatorname{wt}(\operatorname{RM}_{d-1}(m))\leq q\cdot\operatorname{wt}(\operatorname{PRM}_{d}(m-1))=q$. Thus, for $d\leq m(q-1)$ we obtain
@@
$\operatorname{ev}^{\mathbb{A}}_{{m}}(f(1,x_{1},\dots,x_{m}))=\operatorname{ev}^{\mathbb{P}}_{{m}}(f(x_{0},x_{1},\dots,x_{m}))\hskip-2.84544pt\mid_{\{1\}\times\mathbb{F}_{q}^{m}}=(u+v_{\xi,d}),$
where, for $w\in\mathbb{F}_{q}^{q^{m}}$, we denote $w\hskip-2.84544pt\mid_{\{1\}\times\mathbb{F}_{q}^{m}}:=(w_{Q})_{Q\in\{1\}\times\mathbb{F}_{q}^{m}}$.
@@ 3. Recursive decoding algorithm
Fix $m>0$. In this section, we show a decoding algorithm, Algorithm 1, to decode $\operatorname{PRM}_{d}(m)$ if the number of errors is lower than $t_{0}:=\eta_{d}(m)/2$. We denote the corresponding decoder $D^{\mathbb{P}}_{d}(m)$. This algorithm relies on a choice of a decoding algorithm $D^{\mathbb{A}}_{d}(j)$ for RM codes, for each $j\leq m$. To the knowledge of the author, the most efficient decoding algorithm is obtained by considering the Berlekamp-Massey-Sakata algorithm [31, 32] together with the majority voting algorithm [11, 12] to decode RM codes up to their error correction capability [33], or even more depending on the type of errors [7]. A simpler approach could be to just consider majority voting decoding as in [19, 27, 12], which also decodes up to the error correction capability of RM codes, but with higher complexity. With respect to PRM codes, the only specific decoder known is given in [28], in which the authors consider a decomposition of the projective space into affine spaces, while our approach relies on the recursive construction from [34].
@@
As a convention, if at some point we use a decoder whose error correction capability is 0, we assume it returns the received vector if it belongs to the corresponding code, and an error otherwise. With respect to the terminology, by the error correction capability of an $n,k,\delta]$ code $C$ we mean $\left\lfloor\frac{\delta-1}{2}\right\rfloor$, since it can correct any number of errors lower than or equal to $\left\lfloor\frac{\delta-1}{2}\right\rfloor$, or, equivalently, any number of errors lower than $\delta/2$. For completeness, we start by briefly covering the case of PRS codes, then we show how our decoder works for $\operatorname{PRM}_{d}(2)$ to illustrate the main ideas of Algorithm 1, and then we prove that the algorithm works in the general case of $\operatorname{PRM}_{d}(m)$.